OpenStudy (anonymous):

Let f(x) = 5 - x f^{-1}( x ) = How do I solve this??

OpenStudy (gschibby):

You're after the inverse function of f(x)? Just write it out like this y = 5 - x then solve as x(y): x = 5 - y that, I believe is the inverse :)

OpenStudy (gschibby):

So f^-1(x) = 5 - y

OpenStudy (lukecrayonz):

Inverse function means replace every x with y, and everything y with x. So if your equation is y=5-x your inverse is x=5-y Then solve for y and you get the inverse:)

OpenStudy (anonymous):

Ohhh! I see thank you!!!

OpenStudy (lukecrayonz):

http://www.wolframalpha.com/input/?i=inverse+y%3D5-x